OpenStudy (smari306):

Using complete sentences, explain how to find the zeros of the function f(x) = 2x3 – 9x + 3. Please explain? Thanks in advance!

OpenStudy (smari306):

Hello....?

OpenStudy (anonymous):

Let x = (a+b). Then x^3 = (a+b)^3 = a^3 + 3 a b (a + b) + b^3 = a^3 + 3 a b x+ b^3, or x^3 - 3 a b x - (a^3+b^3) = 0 Solve 3 a b = 9, and - (a^3+b^3) = 3 Then the answer is x = a + b x1= 3^(2/3)/(1/2 (-1 + i sqrt(11)))^(1/3) + (3/2 (-1 + i sqrt(11)))^(1/3) x2 = -(((3/2)^(2/3) (1 +i sqrt(3)))/(-1 +i sqrt(11)^(1/3)) - 1/2 (1 -i sqrt(3)) (3/2 (-1 +i sqrt(11)))^(1/3) x3 = -(((3/2)^(2/3) (1 -i sqrt(3)))/(-1 +i sqrt(11))^(1/3)) - 1/2 (1 +i sqrt(3)) (3/2 (-1 + i sqrt(11)))^(1/3)

OpenStudy (whpalmer4):

@dpasingh that has to be adjusted a bit — this equation has 2x^3, not x^3. Your solution is for \[x^3-9x+3=0\]